OpenStudy (anonymous):

the point (-2,-5) lies on the graph of g(x). What point must lie on the graph of g(2x)

OpenStudy (anonymous):

i thought it was (-4,0)

OpenStudy (anonymous):

it only effects the x right?

myininaya (myininaya):

lets do an example k?

myininaya (myininaya):

let's pretend with have a linear equation and lets say (-2,-5) and (0,5) lie on this line. so the slope is {5-(-5)}/{0-(-2)}=10/2=5 so we have y=5x+b and we know a point on this line (0,5) so we have 5=0+b so the y intercept b=5 so the equation of the line through those two points are y=5x+5 so let g(x)=5x+5 g(-2)=5(-2)+5=-10+5=-5 if x=-2 then g(2*(-2))=5(-2)+5=-10+5=-5 so (-4,-5) yes you are right it only affects the x

myininaya (myininaya):

wait oops sorry mistake

myininaya (myininaya):

g(2*(-2))=5(-4)+5=-20+5=-15 (-4,-15)

myininaya (myininaya):

it doesn't only affect the x i;m sorry for saying that in that previous post

myininaya (myininaya):

i have no clue i'm sorry i don't know how to find g(-4) given no other information then g(-2)=-5

OpenStudy (anonymous):

You know that (-2,-5) is on the graph of g(x). So, what about the point g(2(-1)). Plug x=-1 on the function g(2x), you get g(2(-1))=g(-2)=-5. Hence (-1,-5) is a point on the graph of g(2x).

OpenStudy (anonymous):

Another way to see that, is by looking at the function g(2x). You can see that only values of is affected. Set 2x=-2 and solve for x, you get x=-1, which is the same value we got above. S again (-1,-5) is a point on g(2x).

myininaya (myininaya):

omg i'm so stupid

OpenStudy (anonymous):

Haha.. we all have such moments :P